OpenStudy (anonymous):

if the square root of 7x-10 is 12… what is x?

OpenStudy (anonymous):

\[\sqrt{7x-10}=12 \] x=?

OpenStudy (jdoe0001):

\(\bf \sqrt{7x-10} = 12\qquad square \ ^2\textit{ both sides} \\ \quad \\ (\sqrt{7x-10})^2 = (12)^2\implies 7x-10=144\)

OpenStudy (jdoe0001):

then just add 10 to both sides then divide by 7 both